OpenStudy (anonymous):

Find the product and fill in the blanks to write in standard complex number form. (2 + 4i)(7 - 3i) =

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

See, its just simple multiplication. : (2+4i)(7-3i) = 2*7 - 2*3i + 4i*7 - 4i*3i = 14 - 6i + 28i - 12 (i^2) = 14 + 22i - 12(-1) = 14 + 22i + 12 = 26 + 22i

OpenStudy (anonymous):

Thank you. I had 12 as one of the imaginary numbers by mistake.

OpenStudy (anonymous):

Can you help me with one more question?